# Dynamical stability for dense patterns in discrete attractor neural networks

### v3

@@ Introduction.
Figure 1: Storing fixed-points. (a) Color code for combinations of I/O exponent (x-axis) and smoothness (y-axis, log-scale). The activation function for selected combinations is illustrated (insets, top). These combinations are used in b-e, with $\textrm{CV}=2,\theta=-2$. (b,c) The sum (b) and L2-norm (c) of the presynaptic weights of a neuron (i.e. a row of $\mathbf{W}$) as a function of memory load ($\alpha$) for different activation functions (colors as in a). Note the divergence of L2-norm as $\alpha\rightarrow 1$, as predicted by theory. (d) Measures of the weights asymmetry (top) and dynamics non-normality (bottom) at different load values (x-axis), for different activation functions (colors as in a). (e) The fraction of patterns which are correctly stored (thick curves) or stable for recall (thin curves) as a function of memory load ($\alpha$) for different activation functions (colors as in a). (f) Probability density of log-normal distributed patterns at different values of CV (color coded). The inset shows normalized density on a log scale. (g-j) Same as b-e for different values of pattern CV (colors as in f) with $\sigma=1,n=1,\theta=-2$. Solid vs. dotted lines in b,c,e, and g,h,i show numerical vs. theoretical results (with only numerical results shown for stability e,j – see subsequent figures for corresponding theoretical results). Numerical simulations optimized the weights $\mathbf{W}$ according to Eq. 4 at $N=256$.
Network model for auto-associative memory.
@@
While our theoretical results hold for a wide range of activation functions $g\!\left(\cdot\right)$, with additional assumptions detailed below, we consider here the specific case of the soft-rectified power-law to be able to systematically study how stability depends on its parameters (Fig. 1a):
@@
with exponent $n$ and smoothness $\sigma$. The particular form of Eq. 3 is motivated as an approximation to a hard-rectified power-law activation function acting on noisy voltages, where voltage noise is Gaussian with standard deviation $\sigma$: $g\!\left(v\right)=\langle\lfloor v+\sigma z\rfloor^{n}_{+}\rangle_{z\sim{\cal N}\left(0,1\right)}$. Note that due to the noisiness of the activation function (or the equivalent smoothness in Eq. 3), all patterns in the network are dense as firing rates are never exactly zero. We also present results for a noiseless rectified power-law activation function that gives rise to sparse patterns (Details).
@@
Second, unlike previous approaches that considered binary [1, 2] or sparse memory patterns [5, 12], here we specifically focus on dense patterns in which firing rates are never exactly zero: $\mathbf{r}^{\mu}\in\mathbb{\mathbf{R}}_{+}^{N}$. Again, to allow a systematic study of how stability depends on the properties of this distribution, we consider here the specific case of a log-normal distribution of patterns, which we parametrize by its coefficient of variation $\textrm{CV}=\sqrt{\left\langle\delta r^{2}\right\rangle}/\left\langle r\right\rangle$), while fixing its mean at $\left\langle r\right\rangle=1$ (Fig. 1f), where here and in the following $\left\langle\cdot\right\rangle$ and $\delta$ denote averaging across the distribution of memory patterns and computing the deviation from such an average, respectively. For a hard-rectified power law activation function (i.e. $\sigma=0$, Fig. S1b), scaling the mean of memory patterns while keeping their CV constant would cancel in the Jacobian of the dynamics, Eq. 5, thus leaving stability unaffected. For our soft-rectified activation function, this does not hold exactly, but we find the effects on stability to be negligible, Fig. S4. While such patterns are always technically dense, they can approximate sparse distributions with a sufficiently high CV (we call such patterns ‘sparse-like’).
@@
To summarize, we are interested in how the maximal number of patterns $P$ that can be stored as fixed points, and the fraction of these fixed points that are dynamically stable, depend on parameters $\sigma$, $n$, $\theta$, and CV (as well as $f$, the sparseness of patterns, Details). For obtaining this maximum, we optimize $\mathbf{W}$ following Eq. 4.
Mean-field theory for storage capacity.
@@
A replica analysis of the solution for the optimization problem Eq. 4 follows the Gardner approach [11, 12, 18], without assuming specific connection weights as in the Hopfield model [2, 5]. In the Appendix, we derive a mean-field theory for the capacity to store graded, random memory patterns as fixed points of Eq. 1, without assumptions on the weights. The analysis applies to either dense patterns with a strictly monotonic activation function or to sparse patterns with a rectified monotonic activation function. It generalizes previous work in which a rectified-linear activation function and normalized connection weights were assumed [12]. As usual, our analysis assumes $P,N\to\infty$ with a fixed memory load $\alpha=P/N$ and uses the replica technique, assuming the replica symmetry ansatz [19].
@@
In line with previous results [11], our theory predicts that there is a critical capacity $\alpha_{\mathrm{C}}$, such that as $\alpha$ approaches $\alpha_{\mathrm{C}}$, there is no longer a solution to Eq. 4. We find that $\alpha_{\mathrm{C}}$ depends only on pattern sparseness (Details, Fig. S1g) and is independent of other pattern statistics (CV; Fig. 1f), neural thresholds ($\boldsymbol{\theta}$; Eq. 2, Fig. S3a), and activation function details (smoothness $\sigma$ and exponent $n$; Eq. 3, Fig. 1a), thus extending results from neurons with binary [11, 4] or rectified-linear activation functions (corresponding to $\sigma=0$, $n=1$) to our broader class of activation functions ($\sigma\geq 0$, $n>0$) [12]. Results for the sparse case are presented in Fig. S1 (Details). In the dense case, the critical capacity is $\alpha_{\mathrm{C}}=1$ (again, independent from any of the parameters), and theory predicts the following moments for each element of the connection weight matrix:
@@
For a subcritical load, $\alpha=P/N<\alpha_{\mathrm{C}}$, theory can be compared with numerical experiments by directly solving Eq. 4 using off-the-shelf optimizers [20]. The theory provides a good match for both the mean ($N\,\left\langle W\right\rangle$ from Eq. 7; Fig. 1b,g, Fig. S1c, and Fig. S3b) and the L2-norm of a row of the connection weight matrix ($\sqrt{N\,\left\langle W^{2}\right\rangle}$ from Eq. 6; Fig. 1c,h, Fig. S1d, and Fig. S3c). As $\alpha$ approaches $\alpha_{\mathrm{C}}$, the theory predicts the L2-norm of weights to diverge, explaining why there is no solution to Eq. 4 in this regime (Fig. 1e,j, Fig. S1e,f, and Fig. S3e).
To test if the resulting dynamics might coincide with those assumed by the ‘Hebbian’ approach, we measured the weight matrix asymmetry and the deviation from normality of the dynamics around the fixed points (Details). Both are non-negligible and increase with the problem’s load $\alpha$ and pattern variation (Fig. 1d,i, and Fig. S3d), indicating the resulting dynamics are distinct from those of energy-based models.
Importantly, investigating the emergent stability of patterns in the subcritical regime reveals a second phase transition: all patterns tend to be stable up to a new critical value $\alpha_{\mathrm{S}}<\alpha_{\mathrm{C}}$, above which all patterns tend to be unstable (Fig. 1e,j, Fig. S2a,b, and Fig. S3e). In the following, we investigate this phenomenon more closely.
Theory of stability for dense patterns.
For dense patterns, $\alpha_{\mathrm{C}}=1$ and for any $\alpha<\alpha_{\mathrm{C}}$ the solution to Eq. 4 is given in closed-form (Details):
@@
Figure 2: Fixed-point stability. (a) Example spectrum - overlaid eigenvalues of $P$ Jacobians $\mathbf{J_{\mu}}$ of a single problem (parameters indicated by text). Annotations of the theoretical values $\lambda_{\mathrm{bulk}}$ (blue line), $\lambda_{\mathrm{ave}}$ (green line), and $\lambda_{\mathrm{mem}}$ (orange line). The corresponding empirical eigenvalues are marked with matching colors: largest eigenvalue of the bulk (blue dot), the eigenvalue associated with the Jacobian average (green dot), and the eigenvalue associate with the memory pattern (orange dot). Inset shows a near-linear relation between $r^{\mu}_{i}$ (x-axis) and $\phi(r^{\mu}_{i})$ (y-axis), a prerequisite for an outlier $\lambda_{\mathrm{mem}}$. (b) Same as a, for six parameter sets; columns differ by the maximal theoretical value. (c-e) Probability densities at different CV values (color-coded) of the difference between the empirical and predicted value of $\lambda_{\mathrm{bulk}}$ (c), $\lambda_{\mathrm{ave}}$ (d), and $\lambda_{\mathrm{mem}}$ (e), calculated over different choice of the parameters discussed in Figs. 1 and S3, in cases where $\left|\lambda_{\mathrm{bulk}}\right|<10$ (c), $\left|\lambda_{\mathrm{ave}}\right|<10$ (d), $\tau_{\mathrm{mem}}>0.95$ (e).
@@
and build on our previous analysis of the eigenvalue spectrum of random matrices $\mathbf{M}=-\mathbf{I}+\mathbf{X}\,\mathbf{Y}^{\dagger}$ when pairs of the corresponding entries of $\mathbf{X}$ and $\mathbf{Y}$ are (jointly) independent and identically distributed (i.i.d.), correlated Gaussian, and $\alpha$ simply denotes the dimension-ratio (width/height) of $\mathbf{X}$ and $\mathbf{Y}$ (a generalization of $\alpha$ being the load in our case) [22]. This analysis provides an exact result for the support of the eigenvalue spectrum (Fig. S5) and the largest real eigenvalue $\lambda_{\mathrm{bulk}}$ of $\mathbf{M}$ (Details, Eq. 21). Interestingly, we find that, just as the Jacobians of the original network, $\mathbf{J}_{\mu}^{*}$, the Jacobians defined by $\mathbf{M}$ (for any choice of $\mathbf{X}$ and $\mathbf{Y}$) also undergo a phase transition: they are either all stable (for $\alpha<\alpha_{\mathrm{S}}$) or all unstable (for $\alpha>\alpha_{\mathrm{S}}$) (Details, Eq. 22). To apply the general theory of the eigenvalue spectrum of $\mathbf{M}$ to the eigenvalue spectrum of $\mathbf{J}_{\mu}^{*}$ (Eq. 9), we choose $\mathbf{X}$ and $\mathbf{Y}$ such that the variances and covariance of $X_{i\mu}$ and $Y_{i\mu}$ are respectively $c_{\mathrm{ff}}=\langle\delta f\!\left(r,r^{\prime}\right)^{2}\rangle$, $c_{\mathrm{rr}}=\langle\delta r^{2}\rangle$, and $c_{\mathrm{rf}}=\left\langle\delta r\,\delta f\!\left(r,r^{\prime}\right)\right\rangle$, where $f\!\left(r,r^{\prime}\right)=g^{\prime}\!\left(g^{-1}\!\left(r^{\prime}\right)\right)\,\left(g^{-1}\!\left(r\right)+\theta\right)$ for independent variables $r$, $r^{\prime}$. With these substitutions,
@@
where by taking the expectation over $g^{\prime}\!\left(g^{-1}\!\left(\mathbf{r}^{\mu}\right)\right)$, we have again ignored its memory pattern-dependence. Even though the uniform vector is not an eigenvector of $\mathbf{J}_{\mu}^{*}$ in general, when we empirically identify the closest eigenvector (Details), we find that the associated eigenvalue is often close to $\lambda_{\mathrm{ave}}$ (Fig. 2a,d), and can determine the dynamical stability of $\mathbf{J}_{\mu}^{*}$ (Fig. 2b, middle column).
In the second case, we assume that there exists some constant $c$ for which $\mathbf{r}^{\mu}+c\,\boldsymbol{1}$ is an eigenvector of $\mathbf{J}_{\mu}^{*}$. In this case, the following equation must hold (Details):
@@
We first focus on the zero crossing of $\alpha_{\mathrm{S}}$ as a function of the parameters, i.e. the parameter regime where stability is possible at all. For this, we analyze the zero crossing of each of the three terms of Eq. 16 separately by solving the corresponding inequalities: $\alpha_{\mathrm{S}}^{\mathrm{bulk}}>0$, $\lambda_{\mathrm{ave}}<0$, and $\lambda_{\mathrm{mem}}<0$ (assuming $\tau_{\mathrm{mem}}>0.95$). The dependence of each of these zero crossings on $\theta$ is relatively straightforward. The zero crossing of $\alpha_{\mathrm{S}}^{\mathrm{bulk}}$ does not depend on $\theta$ (although the exact value of $\alpha_{\mathrm{S}}^{\mathrm{bulk}}$ when it is greater than zero does, Details). Both $\lambda_{\mathrm{ave}}$ and $\lambda_{\mathrm{mem}}$ depend on $\theta$ linearly (Eq. 12, and Details), and so the condition for their zero crossing can be expressed respectively as
@@
where $d\!\left(r\right)=g^{\prime}\!\left(g^{-1}\!\left(r\right)\right)$, and the RHSs of Eqs. 17 and 18 implicitly depend on the other parameters (Details).
Phase diagrams.
@@
Optimal parameters.
Fig. S4 presents the maximal achievable value for the stability critical load as a function of pattern statistics, as well as the optimal parameters where the optimum is achieved for various choices of $\left<r\right$>. Our theory predicts improvement of stability with pattern variation (Fig. 3c-d, Fig. S4a). Those high-variance memory patterns are sparse-like, with many near-zero values and few large activations, so the improved capacity we report for those echoes the benefit of sparseness in our results (Fig. S2a,e) and previous works [5, 4]. Such sparse-like patterns are also consistent with the known phenomenology of neurons in the brain [23, 24]. As noted above, such patterns require highly asymmetric weights and lead to dynamics distinct from those of energy-based models (Fig. 1i).
@@
In our analysis, the optimal threshold is negative. This is supported by the theory bounding the threshold from above (Fig. 3a-b), and is evident empirically at the optimal parameters for dense patterns (Fig. S4d), as well as for sparse patterns (Fig. S2c). A negative threshold implies that neurons are biased toward activity and are held back by recurrent connections which are dominated by inhibition (i.e. $m<0$), as shown for the near-optimal values $n=1,\sigma=1$ at Fig. 1g. This contrasts with previous proposals for how recurrent neural networks can perform associative memory (a notable exception is the Willshaw network [28]) and with the prevailing view in the field [17], according to which the construction of multiple distinct and stable activity patterns requires global inhibition and selective recurrent excitation.
Discussion.

### v4

@@ Abstract
Figure 1: Storing fixed-points. (a) The activation function for different choices of the threshold $\theta$ (color coded), with I/O exponent $n=1$ and smoothness $\sigma=1$. These combinations are used in b-e, with $\mathrm{CV}=2$. (b,c) The sum (b) and L2-norm (c) of the presynaptic weights of a neuron (i.e. a row of $\mathbf{W}$) as a function of memory load ($\alpha$) for different activation functions (colors as in a). Note the divergence of L2-norm as $\alpha\rightarrow 1$, as predicted by theory. (d) Measures of the weights asymmetry (top) and dynamics non-normality (bottom) at different load values (x-axis), for different activation functions (colors as in a). (e) The fraction of patterns which are correctly stored (thick curves) or stable for recall (thin curves) as a function of memory load ($\alpha$) for different activation functions (colors as in a). (f) Color code for combinations of I/O exponent (x-axis) and smoothness (y-axis, log-scale). The activation function for selected combinations is illustrated (insets, top). (g-j) Same as b-e for combinations of I/O exponent and smoothness (colors as in f) with $\textrm{CV}=2,\theta=-2$. (k) Probability density of log-normal distributed patterns at different values of CV (color coded). The inset shows normalized density on a log scale. (l-o) Same as b-e for different values of pattern CV (colors as in k) with $\sigma=1,n=1,\theta=-2$. Solid vs. dotted lines in b,c,e,g,h,j,l,m,o show numerical vs. theoretical results (with only numerical results shown for stability in e,j,o – see subsequent figures for corresponding theoretical results). Numerical simulations optimized the weights $\mathbf{W}$ according to Eq. 4 at $N=256$.
.2 Network model for auto-associative memory
@@
While our theoretical results hold for a wide range of activation functions $g\!\left(\cdot\right)$, with additional assumptions detailed below, we consider here the specific case of the soft-rectified power-law to be able to systematically study how stability depends on its parameters (Fig. 1a,f):
@@
with exponent $n$ and smoothness $\sigma$. The particular form of Eq. 3 is motivated as an approximation to a hard-rectified power-law activation function acting on noisy voltages, where voltage noise is Gaussian with standard deviation $\sigma$: $g\!\left(v\right)=\langle\lfloor v+\sigma z\rfloor^{n}_{+}\rangle_{z\sim{\cal N}\left(0,1\right)}$. Note that due to the noisiness of the activation function (or the equivalent smoothness in Eq. 3), all patterns in the network are dense as firing rates are never exactly zero. We also present results for a noiseless rectified power-law activation function that gives rise to sparse patterns (Sec. A.1).
@@
Second, unlike previous approaches that considered binary [1, 2] or sparse memory patterns [5, 12], here we specifically focus on dense patterns in which firing rates are never exactly zero: $\mathbf{r}^{\mu}\in\mathbb{\mathbf{R}}_{+}^{N}$. Again, to allow a systematic study of how stability depends on the properties of this distribution, we consider here the specific case of a log-normal distribution of patterns, which we parametrize by its coefficient of variation $\textrm{CV}=\sqrt{\left\langle\delta r^{2}\right\rangle}/\left\langle r\right\rangle$), while fixing its mean at $\left\langle r\right\rangle=1$ (Fig. 1k), where here and in the following $\left\langle\cdot\right\rangle$ and $\delta$ denote averaging across the distribution of memory patterns and computing the deviation from such an average, respectively. For a hard-rectified power law activation function (i.e. $\sigma=0$, Fig. S1b), scaling the mean of memory patterns while keeping their CV constant would cancel in the Jacobian of the dynamics, Eq. 5, thus leaving stability unaffected. For our soft-rectified activation function, this does not hold exactly, but we find the effects on stability to be negligible, Fig. S3. While such patterns are always technically dense, they can approximate sparse distributions with a sufficiently high CV (we call such patterns ‘sparse-like’).
@@
To summarize, we are interested in how the maximal number of patterns $P$ that can be stored as fixed points, and the fraction of these fixed points that are dynamically stable, depend on parameters $\sigma$, $n$, $\theta$, and CV (as well as $f$, the sparseness of patterns, Sec. A.1). For obtaining this maximum, we optimize $\mathbf{W}$ following Eq. 4.
.3 Mean-field theory for storage capacity
@@
A replica analysis of the solution for the optimization problem Eq. 4 follows the Gardner approach [11, 12, 18], without assuming specific connection weights as in the Hopfield model [2, 5]. In Sec. B.1, we derive a mean-field theory for the capacity to store graded, random memory patterns as fixed points of Eq. 1, without assumptions on the weights. The analysis applies to either dense patterns with a strictly monotonic activation function or to sparse patterns with a rectified monotonic activation function. It generalizes previous work in which a rectified-linear activation function and normalized connection weights were assumed [12]. As usual, our analysis assumes $P,N\to\infty$ with a fixed memory load $\alpha=P/N$ and uses the replica technique, assuming the replica symmetry ansatz [19].
@@
In line with previous results [11], our theory predicts that there is a critical capacity $\alpha_{\mathrm{C}}$, such that as $\alpha$ approaches $\alpha_{\mathrm{C}}$, there is no longer a solution to Eq. 4. We find that $\alpha_{\mathrm{C}}$ depends only on pattern sparseness (Sec. A.1, Fig. S1g) and is independent of other pattern statistics (CV; Fig. 1k), neural thresholds ($\boldsymbol{\theta}$; Eq. 2, Fig. 1a), and activation function details (smoothness $\sigma$ and exponent $n$; Eq. 3, Fig. 1f), thus extending results from neurons with binary [11, 4] or rectified-linear activation functions (corresponding to $\sigma=0$, $n=1$) to our broader class of activation functions ($\sigma\geq 0$, $n>0$) [12]. Results for the sparse case are presented in Fig. S1 (Sec. A.1). In the dense case, the critical capacity is $\alpha_{\mathrm{C}}=1$ (again, independent from any of the parameters), and theory predicts the following moments for each element of the connection weight matrix:
@@
For a subcritical load, $\alpha=P/N<\alpha_{\mathrm{C}}$, theory can be compared with numerical experiments by directly solving Eq. 4 using off-the-shelf optimizers [20]. The theory provides a good match for both the mean ($N\,\left\langle W\right\rangle$ from Eq. 7; Fig. 1b,g,l and Fig. S1c) and the L2-norm of a row of the connection weight matrix ($\sqrt{N\,\left\langle W^{2}\right\rangle}$ from Eq. 6; Fig. 1c,h,m and Fig. S1d). As $\alpha$ approaches $\alpha_{\mathrm{C}}$, the theory predicts the L2-norm of weights to diverge, explaining why there is no solution to Eq. 4 in this regime (Fig. 1e,j,o and Fig. S1e,f).
To test if the resulting dynamics might coincide with those assumed by the ‘Hebbian’ approach, we measured the weight matrix asymmetry and the deviation from normality of the dynamics around the fixed points (Sec. A.2). Both are non-negligible and increase with the problem’s load $\alpha$ and pattern variation (Fig. 1d,i,n), indicating the resulting dynamics are distinct from those of energy-based models.
Importantly, investigating the emergent stability of patterns in the subcritical regime reveals a second phase transition: all patterns tend to be stable up to a new critical value $\alpha_{\mathrm{S}}<\alpha_{\mathrm{C}}$, above which all patterns tend to be unstable (Fig. 1e,j,o and Fig. S2a,b). In the following, we investigate this phenomenon more closely.
.4 Theory of stability for dense patterns
For dense patterns, $\alpha_{\mathrm{C}}=1$ and for any $\alpha<\alpha_{\mathrm{C}}$ the solution to Eq. 4 is given in closed-form (Sec. A.3):
@@
Figure 2: Fixed-point stability. (a) Example spectrum - overlaid eigenvalues of $P$ Jacobians $\mathbf{J_{\mu}}$ of a single problem (parameters indicated by text). Annotations of the theoretical values $\lambda_{\mathrm{bulk}}$ (blue line), $\lambda_{\mathrm{ave}}$ (green line), and $\lambda_{\mathrm{mem}}$ (orange line). The corresponding empirical eigenvalues are marked with matching colors: largest eigenvalue of the bulk (blue dot), the eigenvalue associated with the Jacobian average (green dot), and the eigenvalue associate with the memory pattern (orange dot). Inset shows a near-linear relation between $r^{\mu}_{i}$ (x-axis) and $\phi(r^{\mu}_{i})$ (y-axis), a prerequisite for an outlier $\lambda_{\mathrm{mem}}$. (b) Same as a, for six parameter sets; columns differ by the maximal theoretical value. (c-e) Probability densities at different CV values (color-coded) of the difference between the empirical and predicted value of $\lambda_{\mathrm{bulk}}$ (c), $\lambda_{\mathrm{ave}}$ (d), and $\lambda_{\mathrm{mem}}$ (e), calculated over different choice of the parameters discussed in Fig. 1, in cases where $\left|\lambda_{\mathrm{bulk}}\right|<10$ (c), $\left|\lambda_{\mathrm{ave}}\right|<10$ (d), $\tau_{\mathrm{mem}}>0.95$ (e).
@@
and build on our previous analysis of the eigenvalue spectrum of random matrices $\mathbf{M}=-\mathbf{I}+\mathbf{X}\,\mathbf{Y}^{\dagger}$ when pairs of the corresponding entries of $\mathbf{X}$ and $\mathbf{Y}$ are (jointly) independent and identically distributed (i.i.d.), correlated Gaussian, and $\alpha$ simply denotes the dimension-ratio (width/height) of $\mathbf{X}$ and $\mathbf{Y}$ (a generalization of $\alpha$ being the load in our case) [22]. This analysis provides an exact result for the support of the eigenvalue spectrum (Fig. S4) and the largest real eigenvalue $\lambda_{\mathrm{bulk}}$ of $\mathbf{M}$ (Sec. A.4, Eq. 21). Interestingly, we find that, just as the Jacobians of the original network, $\mathbf{J}_{\mu}^{*}$, the Jacobians defined by $\mathbf{M}$ (for any choice of $\mathbf{X}$ and $\mathbf{Y}$) also undergo a phase transition: they are either all stable (for $\alpha<\alpha_{\mathrm{S}}$) or all unstable (for $\alpha>\alpha_{\mathrm{S}}$) (Sec. A.4, Eq. 22). To apply the general theory of the eigenvalue spectrum of $\mathbf{M}$ to the eigenvalue spectrum of $\mathbf{J}_{\mu}^{*}$ (Eq. 9), we choose $\mathbf{X}$ and $\mathbf{Y}$ such that the variances and covariance of $X_{i\mu}$ and $Y_{i\mu}$ are respectively $c_{\mathrm{ff}}=\langle\delta f\!\left(r,r^{\prime}\right)^{2}\rangle$, $c_{\mathrm{rr}}=\langle\delta r^{2}\rangle$, and $c_{\mathrm{rf}}=\left\langle\delta r\,\delta f\!\left(r,r^{\prime}\right)\right\rangle$, where $f\!\left(r,r^{\prime}\right)=g^{\prime}\!\left(g^{-1}\!\left(r^{\prime}\right)\right)\,\left(g^{-1}\!\left(r\right)+\theta\right)$ for independent variables $r$, $r^{\prime}$. With these substitutions,
@@
where by taking the expectation over $g^{\prime}\!\left(g^{-1}\!\left(\mathbf{r}^{\mu}\right)\right)$, we have again ignored its memory pattern-dependence. Even though the uniform vector is not an eigenvector of $\mathbf{J}_{\mu}^{*}$ in general, when we empirically identify the closest eigenvector (Sec. A.6), we find that the associated eigenvalue is often close to $\lambda_{\mathrm{ave}}$ (Fig. 2a,d), and can determine the dynamical stability of $\mathbf{J}_{\mu}^{*}$ (Fig. 2b, middle column).
In the second case, we assume that there exists some constant $c$ for which $\mathbf{r}^{\mu}+c\,\boldsymbol{1}$ is an eigenvector of $\mathbf{J}_{\mu}^{*}$. In this case, the following equation must hold (Sec. A.7):
@@
We first focus on the zero crossing of $\alpha_{\mathrm{S}}$ as a function of the parameters, i.e. the parameter regime where stability is possible at all. For this, we analyze the zero crossing of each of the three terms of Eq. 16 separately by solving the corresponding inequalities: $\alpha_{\mathrm{S}}^{\mathrm{bulk}}>0$, $\lambda_{\mathrm{ave}}<0$, and $\lambda_{\mathrm{mem}}<0$ (assuming $\tau_{\mathrm{mem}}>0.95$). The dependence of each of these zero crossings on $\theta$ is relatively straightforward. The zero crossing of $\alpha_{\mathrm{S}}^{\mathrm{bulk}}$ does not depend on $\theta$ (although the exact value of $\alpha_{\mathrm{S}}^{\mathrm{bulk}}$ when it is greater than zero does, Sec. A.5). Both $\lambda_{\mathrm{ave}}$ and $\lambda_{\mathrm{mem}}$ depend on $\theta$ linearly (Eq. 12, and Sec. A.5), and so the condition for their zero crossing can be expressed respectively as
@@
where $d\!\left(r\right)=g^{\prime}\!\left(g^{-1}\!\left(r\right)\right)$, and the RHSs of Eqs. 17 and 18 implicitly depend on the other parameters (Sec. A.5).
.5 Phase diagrams
@@
.6 Optimal parameters
Fig. S3 presents the maximal achievable value for the stability critical load as a function of pattern statistics, as well as the optimal parameters where the optimum is achieved for various choices of $\left<r\right$>. Our theory predicts improvement of stability with pattern variation (Fig. 3c-d, Fig. S3a). Those high-variance memory patterns are sparse-like, with many near-zero values and few large activations, so the improved capacity we report for those echoes the benefit of sparseness in our results (Fig. S2a,e) and previous works [5, 4]. Such sparse-like patterns are also consistent with the known phenomenology of neurons in the brain [23, 24]. As noted above, such patterns require highly asymmetric weights and lead to dynamics distinct from those of energy-based models (Fig. 1n).
@@
In our analysis, the optimal threshold is negative. This is supported by the theory bounding the threshold from above (Fig. 3a-b), and is evident empirically at the optimal parameters for dense patterns (Fig. S3d), as well as for sparse patterns (Fig. S2c). A negative threshold implies that neurons are biased toward activity and are held back by recurrent connections which are dominated by inhibition (i.e. $m<0$), as shown for the near-optimal values $n=1,\sigma=1$ in Fig. 1l. A negative threshold was already reported as a condition for stability in some other, unrelated models [28, 29, 30]. Note that the inhibition-dominated connections in our model differ from previous proposals for how recurrent neural networks can perform associative memory and with the prevailing view in the field [17], according to which the construction of multiple distinct and stable activity patterns requires global inhibition and selective recurrent excitation.
.7 Discussion